OpenStudy (anonymous):

Somebody help me answer this question I will Medal and fan them. Jamison’s father says he will put some money in Jamison’s bank account for his education. The amount he will deposit will either be the number of quarters laid end to end that are equal to Jamison’s height or the number of nickels stacked on top of each other that are equal to Jamison’s height. Jamison wants the most money possible. Each nickel is about --1 12 inch thick. Each quarter is about 1 inch wide. Jamison’s height is 4 feet, 8 inches. (a) Without calculating the total amount he would receive. predict which deposit option will give Jamison the most money.

OpenStudy (luffingsails):

First step is to get the same units. So, take Jamison's height all in inches. Can you do that?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

There will be 12 nickels...

OpenStudy (luffingsails):

Then once you have done that, you can divide his height in inches by the height of the nickel and the width of the quarter.

OpenStudy (luffingsails):

Actually, if the quarter is 1/12 of an inch there will be a lot more quarters than that.

OpenStudy (luffingsails):

4 feet and 8 inches = 4 * 12 + 8 = 56 inches. 56 / (1/12) = 56 * 12 = 672